Name the property of real numbers illustrated below. Be specific with the naming (include “of addition” or “of multiplication,” if applicable).
8 * (4 * 9) = (8 * 4) * 9


Solve for x:
(x – 5)/3 – 3/2 = (4x + 7)/6

Answer by Edwin McCravy(20086)


Name the property of real numbers illustrated below. Be specific with the
naming (include “of addition” or “of multiplication,” if applicable).

8 * (4 * 9) = (8 * 4) * 9 

Notice that on the left side, the parentheses around the 4 * 9 
ASSOCIATES the MULTIPLICATION of the 4 times the 9, and on the 
right side, the parentheses around the 8 * 4 on the left ASSOCIATES the
MULTIPLICATION of the 8 times the 4.  Therefore this property is
called the ASSOCIATIVE property of MULTIPLICATION.


Solve for x:
      (x – 5)/3 – 3/2 = (4x + 7)/6

Multiply every term on the left and right by the LCD of 6.

(6)(x - 5)/3 - (6)3/2 = (6)(4x + 7)/6

Cancel the denominator 3 into the (6) in the first term on the left.
Cancel the denominator 2 into the (6) in the second term on the left.
Cancel the denominator 6 into the (6) in the term on the right.

 2              3        1
(6)(x - 5)/3 - (6)3/2 = (6)(4x + 7)/6
           1        1               1

    (2)(x - 5) - (3)3 = (4x + 7)

         2(x - 5) - 9 = 4x + 7
 
Now you can finish and get x = -13

Edwin